OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P? (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 64 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 8

OpenStudy (anonymous):

Can anyone help? Please i'd truly appreciate it

OpenStudy (anonymous):

(x – 2)^2 + (y + 1)^2 = 64 . Plug in the values into the equation of the circle.

OpenStudy (anonymous):

yeah sorry i'm solving it. I didn't notice you responded lol

OpenStudy (anonymous):

Doesn't seem like that's the answer.....

OpenStudy (anonymous):

Would it be the last one?

OpenStudy (anonymous):

Nope, It is definitely (x – 2)^2 + (y + 1)^2 = 64

OpenStudy (cwrw238):

equation of a circle is (x - a)^2 + (y - b)^2 = r^2 where center is (a,b) and radius = r